# Accelerating AllReduce with a Persistent Straggler

### v1

@@ 13            $Q_{r}\leftarrow$ available ranks holding other active chunks in $A\setminus\{\min_{c}A\}$
16                   Match rank $g\leftrightarrow h$; $g$ sends $c$; $h$ sends $c^{\prime}$; mark $g$ and $h$ as unavailable
@@
19      Match $p_{i}\leftrightarrow q_{i}$ for $p_{i}\in P_{r},q_{i}\in Q_{r}$; exchange their active chunks
@@ 20      Append a list of round $r$’s matchings to schedule
              // Add newly active chunk to $A$
@@ 3.1 Algorithm Design
We consider a setting with $n$ ranks, $10,{\ldots},n{-}$ 1. Each rank has a data buffer of $s$ bytes that is evenly divided into $n{-}1$ chunks, each of size $\frac{s}{n-1}$ and indexed as $c_{0},{\ldots},c_{n-2}$. Without loss of generality, we assume rank $n{-}1$ is the persistent straggler; however, the rank of the actual persistent straggler can simply be swapped with rank $n{-}1$ in the schedules computed by StragglAR.
Precondition and postcondition. StragglAR uses a precondition where the $n{-}1$ non-straggler ranks have completed a ReduceScatter among themselves, ideally fully overlapping with the delay of the straggler. This is likely, since the ReduceScatter is 2×2$\times$2 × faster than AllReduce and is efficiently implemented by libraries like NCCL or MPI [19]. This precondition reduces subsequent communication overhead with the straggler (example in Fig. 4(a)). We refer to the reduced chunks after this step as partially reduced because they contain all but the straggler’s data. StragglAR computes an optimized communication schedule to transition from this precondition to the standard AllReduce postcondition, where all ranks, including the straggler, possess a fully reduced buffer.
The communication schedule consists of multiple rounds, each represented by a set of matchings, similar to algorithms for bandwidth-optimal broadcast [27]. In each round, matchings specify the pairs of ranks that communicate and which data chunks they exchange simultaneously. Each rank participates in exactly one matching per round. Since we fix the chunk size per round, the only lever to minimize communication time is to reduce the number of rounds. At the very least, $n{-}1$ rounds are required to disseminate all $n{-}1$ chunks. StragglAR can complete the AllReduce in $n{+}\log n{-}2$ rounds, yielding a $\beta$ cost of $\tfrac{n+\log n-2}{n-1}s\beta$, which is lower than existing algorithms (see §3.2). Algorithm 1 summarizes the schedule computation when $n$ is a power-of-two, a common case in practice and consistent with canonical collective algorithms [28]. We discuss extending this to even, non-power-of-two cluster sizes in §C. Fig. 4(a) visualizes the StragglAR schedule for $n{=}4$.
Pairing with the straggler. In round $r$, rank $r$ pairs with the straggler (rank $n{-}1$) to fully reduce chunk $c_{r}$. Thus, after $n{-}1$ rounds, the straggler has fully reduced its entire buffer and for each chunk $c_{i}$, at least one non-straggler GPU (certainly GPU $i$) holds its fully reduced copy. Due to this invariant, $c_{r}$ can only be sent to other GPUs from round $r{+}1$ onwards, since it is only fully reduced in round $r$.
Phase 1: Everyone gets a chunk! In the first $\log n$ rounds, the goal is to ensure that every rank obtains a fully reduced chunk. The algorithm guarantees this via the following two rules:
  • In round $r\in[1,\log n{-}1]$, rank $r{-}1$ sends chunk $c_{r{-}1}$ to rank $r{-}1{+}\log n$.
@@
Phase 2: Careful doubling of active chunks. At the beginning of round $r=\log n$, every rank has exactly one fully reduced chunk. Specifically, exactly one rank, $r{-}1$, has $c_{r-1}$. Two ranks have $c_{r-2}$ because it was first fully reduced in round $r{-}2$ and then transmitted to another rank in round $r{-}1$. By the same pattern, four GPUs have $c_{r-3}$, etc.
Definition 3.0.1.
@@
An active chunk $c_{j}$ is a chunk that has been fully reduced with the straggler by round $r>j$, but has yet to propagate to all $n$ ranks.
(a) StragglAR schedule for a 4-GPU setup.
(b) Example of matching process with active chunks.
Figure 4: Algorithm design. (b) shows the matching process with 8 GPUs in round $r{=}4$. Ranks 3 and 5 cannot match because 5 is in the critical window and 3’s active chunk was only recently unlocked.
@@
As we prove in §B, every non-straggler rank possesses exactly one active chunk at any point in time. StragglAR aims to propagate active chunks to all ranks as quickly as possible. Being cautiously greedy works because maximizing propagation (i.e., doubling) in the current round frees up ranks to help propagate other chunks that are fully reduced later on, in subsequent rounds. Intuitively, an active chunk requires a minimum of $\log n$ rounds to propagate fully, by doubling in each round. This implies that an active chunk $c_{j}$ is due for full propagation by round $j+\log n$.
@@
Algorithm 1 enables every active chunk to double in every round by ensuring that every rank transmits its active chunk and receives a different active chunk. This still ensures that every rank only has one active chunk at any point in time because one of those active chunks will expire in the next round. We represent the problem in each round as a bipartite matching problem; luckily, we can construct the matching intuitively without a separate algorithm. In each round $r$, we define a graph $G_{r}$ with vertices $V$, corresponding to non-straggler ranks. We can characterize each rank $v\in V$ by its present active chunk. We partition the ranks into two disjoint sets: $P_{r}$, ranks $v\in V$ whose active chunk in round $r$ is $c_{r{-}\log{n}}$ (i.e., the oldest active chunk that must propagate fully by this round to meet its $\log n$ propagation deadline), and $Q_{r}$, the remaining non-straggler ranks in $V$. By the doubling property, it follows that $|P_{r}|=2^{\log{n}{-}1}=\tfrac{n}{2}$ and thus $|Q_{r}|=n{-}1{-}|P_{r}|=\tfrac{n}{2}{-}1$.
@@
This allows us to define an invariant: For every round $r\in[\log n,n{-}2]$, rank $r\in P_{r}$. The invariant ensures that ranks receiving a new active chunk have already received the active chunk due in this round ($c_{r{-}\log n}$) so that it fully propagates within its $\log n$ round deadline. In other words, the number of active chunks must stay bounded at $\log n$; a rank cannot introduce a new active chunk (by pairing with the straggler) without another active chunk expiring. In round $r$, we restrict the matching algorithm to exclude rank $r$ since it pairs with the straggler. This ensures $|P_{r}|=|Q_{r}|=\frac{n}{2}{-}1$ for that round. Next, we show how we can construct the matchings and ensure $r\in P_{r}$.
@@
Now, we have defined two equally sized, disjoint sets of ranks, $P_{r}\setminus\{r\}$ and $Q_{r}$. However, we cannot generate arbitrary matchings, even though $u$ and $v$ lack each other’s active chunks, because this could violate the invariant later on. For example, rank $r{+}1$ should not receive a chunk that will remain active in round $r{+}1$, as rank $r{+}1$ must pair with the straggler in round $r{+}1$ and thus cannot help double any active chunks. Specifically, rank $r{+}1$ can only receive a chunk $c_{j},j\leq r{+}1{-}\log{n}$, rank $r+2$ can only receive $c_{j},j\leq r{+}2{-}\log n$, etc. Because ranks within this $\log n$-sized critical window (i.e., ranks to pair with the straggler in the next $\log n$ rounds) are more constrained, we match these ranks $j\in[r{+}1,r{+}\log(n)]$ first, by selecting any unmatched neighbor outside the critical window that has the oldest available active chunk. After finalizing these matchings, the remaining ranks can be matched arbitrarily since, by definition, every $u\in P_{r}$ and $v\in Q_{r}$ lack each other’s active chunks.
@@
After $n{-}1$ rounds, every chunk has been fully reduced, and the straggler has fully reduced its buffer. To fully propagate remaining chunks quickly, the straggler can be paired with arbitrary ranks, but always sends the final chunk. (This enables the straggler to be added to $Q_{r}$ so that $|P_{r}|=|Q_{r}|=\tfrac{n}{2}$.)
@@ 3.2 Communication Complexity
We can guarantee the number of rounds in schedules computed by StragglAR when $n$ is a power of 2.
@@
We provide the proof in §B. The intuition is that StragglAR ensures every chunk $c_{r}$ that is fully reduced in round $r$ fully propagates within $\log n$ rounds. A special case is the final chunk $c_{n-1}$, which propagates in $\log n{-}1$ rounds after round $n{-}1$ (when it is fully reduced) due to the straggler now being available to help with transmission, thereby achieving $n{+}\log n{-}2$ total rounds. Ensuring perfect doubling for each chunk in every round requires matching ranks optimally for the current round without compromising on propagation in future rounds. This demands special care due to the pre-determined matching (for symmetry breaking) between rank $r$ and the straggler in round $r$.
$\alpha$-$\beta$ cost. For analysis, we assume that the persistent straggler is delayed long enough to mask ReduceScatter precondition (supported by findings in §2). Since StragglAR achieves $n{+}\log n{-}2$ rounds with $\tfrac{s}{n-1}$ (i.e., chunk size) bytes transmitted parallelly per round, the total time taken is
$T_{SAR}=(n{+}\log{n}{-}2)\cdot\alpha+\tfrac{n{+}\log n{-}2}{n{-}1}s\cdot\beta.$
As shown in Table 1, the $\beta$ cost of StragglAR is substantially lower than that of baselines for power-of-2 $n\geq 4$. Moreover, StragglAR’s advantage over baselines only grows as $n$, the size of the GPU cluster, scales. The scaling behavior of the baseline $\beta$ cost is $\lim_{n\to\infty}\tfrac{2(n{-}1)}{n}=2$ whereas StragglAR achieves $\lim_{n\to\infty}\tfrac{n{+}\log n{-}2}{n{-}1}=1$, suggesting 2×2$\times$2 × faster AllReduce with many GPUs and a persistent straggler. For $\alpha$ cost, which is typically negligible for larger buffers, StragglAR scales better than Ring but more poorly than Recursive Halving/Doubling. We note that even our offline algorithm is fast: StragglAR computes the schedule for a 256-GPU cluster in 1.04 seconds.
Algorithm Latency cost Bandwidth cost
Ring $2(n{-}1)\cdot\alpha$ $\tfrac{2(n-1)}{n}\cdot s\cdot\beta$
RH/D $2\log n\cdot\alpha$ $\tfrac{2(n-1)}{n}\cdot s\cdot\beta$
StragglAR $\left(n{+}\log{n}{-}2\right)\cdot\alpha$ $\tfrac{n{+}\log n{-}2}{n{-}1}\cdot s\cdot\beta$
Table 1: Communication complexity of different AllReduce algorithms for power-of-2 values of $n$, assuming a sufficient straggler delay.
@@ 4 Experiments
We implement the algorithm to compute schedules and a runtime to execute these schedules in multi-GPU setups. Standard profiling tools can be run beforehand to diagnose persistent stragglers [22, 1, 3].
@@
Schedule generator. We implement the StragglAR algorithm in Python to generate AllReduce schedules. The algorithm takes $n$, the number of GPUs, as input and outputs a list for the schedule. Each element in this list is the set of matchings for that round, where each matching specifies which GPUs communicate and which chunks to send/receive. StragglAR is run offline and only needs to be run once for a value of $n$. If the straggler is not rank $n{-}1$, it can simply be swapped with $n{-}1$ offline.
@@
Runtime. We implement the StragglAR schedules in CUDA using the NCCL Point-to-Point API (ncclSend() and ncclRecv()[29, 20] for communication and separate kernels for reduction. After completing ncclReduceScatter() among non-straggler GPUs, we invoke the CUDA kernel to execute the StragglAR schedule. There are two barriers: (1) among the non-straggler GPUs for ReduceScatter, and (2) among all GPUs, before launching the schedule. Both ReduceScatter and StragglAR are packaged into a CUDA API that has a similar functionality to ncclAllReduce().
4.1 Baselines
Given the switched connectivity of GPU scale-up domains, we compare StragglAR to the hardest bandwidth-efficient AllReduce algorithms. For fair comparison of the algorithmic contribution, we implement baselines using the NCCL P2P API and the same CUDA compute kernels as StragglAR.
Ring [30]. This is the standard bandwidth-efficient algorithm implemented by NCCL for AllReduce. It divides the buffer into chunks of size $\tfrac{s}{n}$ and consists of $2(n-1)$ rounds in a ring-like communication pattern. The total runtime is $T_{Ring}=2(n-1)\cdot\alpha+2\tfrac{n-1}{n}s\cdot\beta$.
Recursive halving/doubling (RHD) [28]. Also called the Butterfly algorithm, this algorithm—like Ring—consists of two phases. In the first phase, the buffer is first divided in $1/2$ with one partner, then in $1/4$ with another partner, etc. to achieve a ReduceScatter. Then, a mirror-image AllGather completes to propagate all fully reduced chunks. The total runtime is $T_{RHD}=2\log n\cdot\alpha{+}2\tfrac{n-1}{n}s\cdot\beta$.
Broadcast. This is a straggler-aware baseline that assumes the non-stragglers complete an AllReduce (2×2$\times$2 × the time as ReduceScatter) during the straggler delay. After the straggler is ready, it exchanges its entire buffer with any other rank to fully reduce the entire buffer and then initiates a broadcast with $\log n$ rounds and $s$ bytes per round. The total runtime is $T_{Bcast}=\log n\cdot\alpha{+}\log n\cdot s\cdot\beta$.
4.2 Evaluating StragglAR on Multi-GPU Servers

### v2

@@ 13    $Q_{r}\leftarrow$ available ranks holding other active chunks in $A\setminus\{\min_{c}A\}$
16       Match rank $\leftrightarrow h; gg$ sends $c$; $h$ sends $c^{\prime}$; mark $g$ and $h$ as unavailable
@@
19 Match $p_{i}\leftrightarrow q_{i}$ for $p_{i}\in P_{r},q_{i}\in Q_{r}$ to exchange their active chunks
@@ 20  Append a list of round $r$’s matchings to schedule
    // Add newly active chunk to AA
@@ 3.1 Algorithm Design
We consider a setting with $n$ ranks, $0,{\ldots},n{-}1$. Each rank has a buffer of $s$ bytes that is evenly divided into n1n{-}1 chunks, each of size $\frac{s}{n-1}$ and indexed as $c_{0},{\ldots},c_{n-2}$. Without loss of generality, we assume rank n1n{-}1 is the straggler in this section; however, by symmetry, the algorithm applies regardless of which rank is the straggler (as further shown by our evaluation in §4).
Precondition and postcondition. StragglAR uses a precondition where the n1n{-}1 non-straggler ranks have completed a ReduceScatter among themselves, ideally overlapped with the delay of the straggler. This is feasible, since ReduceScatter is inherently $2\times$ faster than AllReduce and is efficiently implemented by CCLs [Thakur et al., 2005]. We refer to the reduced chunks after this step as partially reduced because they contain all but the straggler’s data. StragglAR computes an optimized communication schedule to transition from this precondition to the standard AllReduce postcondition, where all ranks, including the straggler, possess a fully reduced buffer.
The communication schedule consists of multiple rounds, each represented by a set of matchings, similar to some broadcast algorithms [Bar-Noy et al., 2000]. In each round, matchings specify the pairs of ranks that communicate and which data chunks they exchange simultaneously. Each rank participates in exactly one matching per round. Since we fix the chunk size per round, the only lever to minimize communication time is to reduce the number of rounds. At the very least, n1n{-}1 rounds are required to transmit all n1n{-}1 chunks. StragglAR completes the AllReduce in $n{+}\log n{-}2$ rounds, yielding a $\beta$ cost of $\tfrac{n+\log n-2}{n-1}s\beta$, which is less than the known lower bound of $2\tfrac{n-1}{n}s\beta$ (see §3.2). Alg. 1 summarizes the schedule computation when $n$ is a power-of-two, with modifications for non-power-of-two cluster sizes in §D. Fig. 4(a) visualizes the StragglAR schedule for n=4n{=}4.
Straggler pairings. In round $r$, rank $r$ pairs with the straggler to fully reduce $c_{r}$. After n1n{-}1 rounds, the straggler’s buffer is fully reduced and each chunk $c_{i}$ has been fully reduced on at least one non-straggler rank. Hence, $c_{r}$ can propagate only from round r+1r{+}1 onward, once fully reduced.
Phase 1. In the first $\log n$ rounds, every rank obtains a fully reduced chunk, guaranteed by two rules:
  • In round $r\in[1,\log n{-}1]$, rank r1r{-}1 sends chunk $c_{r{-}1}$ to rank $r{-}1{+}\log n$.
@@
Phase 2. At the start of round $r=\log n$, every rank has exactly one fully reduced chunk. Exactly one rank, r1r{-}1, has $c_{r-1}$. Two ranks have $c_{r-2}$ because it was first fully reduced in round r2r{-}2 and then transmitted to another rank in round r1r{-}1. By the same pattern, four ranks have $c_{r-3}$, etc.
@@
Definition. An active chunk $c_{j}$ is a chunk that has been fully reduced with the straggler by round $r>j$, but has yet to propagate to all $n$ ranks.
(a) Example StragglAR schedule for 4 GPUs.
(b) Example matching process, round r=4r{=}4 with 8 GPUs.
Figure 4: Algorithm design (straggler in red). In (b), ranks 3 and 5 cannot match: 5 is in the critical window and 3’s chunk was just unlocked. Pairing them would prevent 3’s chunk from doubling.
@@
As we prove in §C, every non-straggler possesses exactly one active chunk at any point in time. StragglAR aims to propagate active chunks to all ranks as quickly as possible. Each active chunk requires a minimum of $\log n$ rounds to propagate fully, by doubling in each round, so an active chunk $c_{j}$ is due for full propagation to all ranks by round $j+\log n$.
@@
Algorithm 1 enables every active chunk to double in every round by ensuring that every rank transmits its active chunk and receives a different active chunk, with the oldest active chunk expiring in each round. We model each round as a bipartite matching problem between two disjoint sets of non-straggler ranks: $P_{r}$, ranks whose active chunk in round $r$ is $c_{r{-}\log{n}}$ and $Q_{r}$, the others. By the doubling property, it follows that $|P_{r}|=2^{\log{n}{-}1}=\tfrac{n}{2}$ and thus $|Q_{r}|=n{-}1{-}|P_{r}|=\tfrac{n}{2}{-}1$.
@@
This allows us to define an invariant: For every round $r\in[\log n,n{-}2]$, rank $r\in P_{r}$. This guarantees the $\log n$ propagation deadline, by ensuring that ranks receiving a new active chunk (by pairing with the straggler) have already received the chunk due in that round ($c_{r{-}\log n}$). Thus, we exclude rank $r$ from the matching process since it pairs with the straggler, enabling $|P_{r}|=|Q_{r}|=\frac{n}{2}{-}1$.
@@
Now, we have two equal-sized, disjoint sets of ranks, $P_{r}\setminus\{r\}$ and $Q_{r}$. However, arbitrary matchings between the two sets risk violating the invariant in the future. In particular, ranks in the next $\log n$ rounds (the critical window) cannot receive chunks that would remain active when they must pair with the straggler (Fig. 4(b)). For example, rank r+1r{+}1 may only receive $c_{j}$ with $j\leq r{+}1{-}\log n$. To enforce this, we first match ranks $j\in[r{+}1,r{+}\log{n}]$ in the critical window with partners outside this window that have the oldest available active chunk. After finalizing these matchings, the remaining ranks can be paired arbitrarily since every $u\in P_{r}$ and $v\in Q_{r}$ lack each other’s active chunks.
@@
After n1n{-}1 rounds, every chunk has been fully reduced, and the straggler has fully reduced its buffer. To fully propagate remaining chunks quickly, the straggler can be paired with arbitrary ranks, but always sends the final chunk. (This enables the straggler to be added to $Q_{r}$ so that $|P_{r}|=|Q_{r}|=\tfrac{n}{2}$.)
@@ 3.2 Communication Complexity
We provide the proof in §C. Intuitively, each chunk $c_{r}$ propagates within $\log n$ rounds of being fully reduced. The final chunk $c_{n-1}$ only requires $\log n{-}1$ rounds to propagate, as the straggler can also help transmit it, thereby achieving $n{+}\log n{-}2$ total rounds. Achieving this bound requires matching ranks optimally for the current round without compromising on future propagation, where special care must be taken due to the pre-determined straggler pairings.
Algorithm Latency Bandwidth
Ring $2(n{-}1)\alpha$ $\tfrac{2(n-1)}{n}s\beta$
RHD $\boldsymbol{2(\log n)\alpha}$ $\tfrac{2(n-1)}{n}s\beta$
StragglAR $(n{+}\log n{-}2)\alpha$ $\boldsymbol{\tfrac{n{+}\log n{-}2}{n{-}1}s\beta}$
Table 1: AllReduce complexity with straggler delays.
When the straggler is delayed long enough to mask the ReduceScatter precondition (supported by Fig. 2(a)), StragglAR’s exposed communication time consists of $n{+}\log n{-}2$ rounds with $\tfrac{s}{n-1}$ bytes sent per round. Thus, the total time taken is
$T_{\text{SAR}}=(n+\log n-2)\,\alpha+\frac{n+\log n-2}{n-1}\,s\beta.$
Under these conditions, StragglAR achieves much lower $\beta$ cost than today’s known bandwidth-optimal lower bound (see Table 1), and this advantage only grows as $n$ (number of GPUs) scales. The scaling behavior of known $\beta$-optimal algorithms is $\lim_{n\to\infty}\tfrac{2(n{-}1)}{n}s\beta=2s\beta$, whereas StragglAR achieves $\lim_{n\to\infty}\tfrac{n{+}\log n{-}2}{n{-}1}s\beta=s\beta$, implying $2\times$ speedups in large-scale settings with a straggler. Even in the worst case, where none of the ReduceScatter precondition can be overlapped (e.g., no stragglers), StragglAR remains competitive: since a ReduceScatter with n1n{-}1 ranks only incurs $\tfrac{n{-}2}{n{-}1}s\beta$ cost, StragglAR achieves $\lim_{n\to\infty}\tfrac{n{-}2}{n{-}1}s\beta+\tfrac{n{+}\log n{-}2}{n{-}1}s\beta=2s\beta$. While $\alpha$ cost is typically negligible for large buffers, StragglAR scales better in $\alpha$ cost than Ring but more poorly than RHD.
@@ 4 Experiments
We implement the StragglAR algorithm to compute schedules, as well as a CUDA runtime that executes these schedules in multi-GPU setups.
Detecting stragglers. One advantage of StragglAR is that its initial ReduceScatter can be eagerly executed as soon as the first n1n{-}1 ranks are ready. If there is sufficient delay between the slowest rank and these n1n{-}1 ranks, StragglAR achieves its ideal performance; otherwise, performance falls in the range between the ideal and worst-case bounds discussed in §3 and §4.3. To consistently support ideal-case performance, the runtime can leverage online straggler detection tools [Zhao et al., 2024a] and pass the straggler’s rank to the backend for conditional execution of the appropriate schedule. In our end-to-end experiments (§4.2), we fix the rank that StragglAR assumes to be the straggler (by profiling the workload ahead of time and picking a likely straggler rank, see §4.2). This stress-tests StragglAR, as there are many iterations in which the algorithm encounters its worst-case performance (i.e., no precondition overlap), when a different rank is the straggler or there is no straggler at all.
@@
Schedule generator. We implement StragglAR in Python to generate AllReduce schedules. The algorithm takes the number of GPUs as input and outputs a schedule with matchings, specifying for each round which GPUs communicate and which chunks to send/receive. Schedule generation is run once offline and is fast—StragglAR computes the schedule for a 256-GPU cluster in <1.04 seconds.
@@
Runtime. We implement the StragglAR schedules using the NCCL Point-to-Point (P2P) API [NCCL, ] (also used in prior work [Hwang et al., 2023]), for GPU communication, combined with custom CUDA kernels for reduction. After completing ncclReduceScatter() among non-straggler GPUs, we invoke the runtime to execute the StragglAR schedule. We package both ReduceScatter and StragglAR into an API that has the same functionality as ncclAllReduce().
Baselines. Given the homogeneous, switched connectivity of modern scale-up networks, we compare StragglAR to the strongest bandwidth-optimal AllReduce algorithms for this setting. Our baselines are (1) Ring [Patarasuk and Yuan, 2009, NVIDIA, 2024a] and (2) Recursive Halving/Doubling (RHD) [Bruck et al., 1994], which are the gold-standard bandwidth-optimal algorithms in homogeneous networks [Hu et al., 2025]; (3) MSCCL [Cowan et al., 2023], the only recent work we are aware of to synthesize a new algorithm for scale-up domains; and (4) Broadcast, a naive straggler-aware baseline. The MSCCL baseline uses their allpairs algorithm: a one-round ReduceScatter where each GPU splits its bandwidth across all peers, followed by a one-round mirror AllGather. This achieves the same $2s\beta$ bandwidth cost as classical algorithms but reduces latency to $2\alpha$ on switched networks. The Broadcast baseline is straggler-aware, as non-stragglers complete an AllReduce during the straggler delay and then finish with a pairwise-exchange broadcast once the straggler arrives. More details on baselines can be found in §E. For fair comparison of the algorithmic contribution, we implement baselines using the NCCL P2P API and the same CUDA compute kernels as StragglAR.
4.1 Benchmarking StragglAR on Multi-GPU Scale-Up Domains